# Improving absolute separability bounds for arbitrary dimensions

### v1

@@ 4.1 Symmetric absolute separability and absolute PPT
The proof is analogous to that of Proposition 1. ∎
When it comes to comparing our criterion in Eq. (20) with existing SAS criteria [22, 44, 45], we find the extremal states detected by our conditions not captured by any of these. Thus, our condition provides a broader set of states detected. Nevertheless, our proof considers only SAP, for which we have not found analytical expressions in the literature. The sets of (S)AS and (S)AP being equivalent remains an open problem to possibly consider our criteria for separability or to disprove the conjectured equality of both sets [24].
The negative bound on $\alpha_{-}$ in Theorem 9, as already discussed, is not tight in general, and we foresee a dependence on $d$ for it. From [19], it is possible to compute the exact bound for the symmetric bipartite case, when $N=2$ and the local dimension $d$ is arbitrary. For the specific case of a symmetric $2-$qutrit system, we find the following results, which again only depend on a single eigenvalue of the density matrix.
Theorem 10.
Let $\rho\in\mathcal{S}(\mathbb{C}^{3}\otimes\mathbb{C}^{3})$ be a symmetric state of a $2-$qutrit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{7}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{3}{16},$ (21)
then, $\rho$ is symmetric absolutely PPT.
Proof.
The positive and negative bounds on $\alpha\in[-2/3,1]$ can be computed by substituting the vector of eigenvalues $\left(1-(6-1)/(6+\alpha_{\pm},1/(6+\alpha_{\pm}),\cdots,1/(6+\alpha_{\pm}))\right)$ on Corollary 3 of [19]. The remainder of the proof is analogous to the one of Theorem 3. ∎
From the two simplexes given by the minimal and maximal eigenvalue criteria, we derive the following sufficient condition, as done previously for similar cases.
Proposition 6.
Let $\rho\in\mathcal{S}(\mathbb{C}^{3}\otimes\mathbb{C}^{3})$ be a symmetric state of a $2-$qutrit system with $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,5}$ its corresponding eigenvalues in non-decreasing order. If
$5\lambda_{0}+5\lambda_{1}+4\lambda_{2}\geq 2,$ (22)
then $\rho$ is symmetric absolutely PPT.
Proof.
The proof is analogous to that of Proposition 3. ∎
For higher dimensional bipartite systems $d>3$, one may recover a tight value of negative $\alpha$ by relaxing the LMIs presented in [19]. However, the number of such LMIs to consider prohibits a scalable characterization beyond $d=4$. For this last case, we derive the bound $\alpha_{-}=-0.625$ from their Conjecture 1. Strong numerical evidence suggests that such bound is indeed tight. Thus providing the following criteria for two ququart symmetric states.
Theorem 11.
Let $\rho\in\mathcal{S}(\mathbb{C}^{4}\otimes\mathbb{C}^{4})$ be a symmetric state of a $2-$ququart system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{11}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{8}{75},$ (23)
then, $\rho$ is symmetric absolutely PPT.
Proof.
The positive and negative bounds on $\alpha\in[-5/8,1]$ can be computed by substituting the vector of eigenvalues $\left(1-(10-1)/(10+\alpha_{\pm},1/(10+\alpha_{\pm}),\cdots,1/(10+\alpha_{\pm}))\right)$ on Conjecture 1 of [19]. We have gathered numerical evidence that the value of $\alpha_{-}$ is, in fact, tight, supporting the conjecture. The rest of the proof is analogous to the one of Theorem 3. ∎
From the two simplexes given by the minimal and maximal eigenvalue conditions, we derive the following sufficient condition, as done previously for similar cases.
Proposition 7.
Let $\rho\in\mathcal{S}(\mathbb{C}^{4}\otimes\mathbb{C}^{4})$ be a symmetric state of a $2-$ququart system with $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,9}$ its corresponding eigenvalues in non-decreasing order. If
$3\lambda_{4}+13\cdot\sum_{i=0}^{3}\lambda_{i}\geq 5,$ (24)
then $\rho$ is symmetric absolutely PPT.
Proof.
The proof is analogous to that of Proposition 3. ∎
To close the section, we summarize the main criteria established here for SAS. We first provide tight SAS criteria for two and three qubits. Then, we examine the twin problem of SAP and present new non-tight criteria for arbitrary dimensions and tight ones for two qutrits and ququarts systems. For these, we have strong numerical evidence for the tightness of our bounds. Finally, it would be beneficial to develop scalable criteria for SAP and study its interplay with SAS, which we leave for subsequent investigations.
@@ 5 Conclusions
In this work we have presented a new method to tackle the problem of absolute separability (AS) and absolute PPT (AP) based on inverting non completely positive maps. The AS (AP) sets are defined as those formed by the states whose separability (positivity under partial transposition) is preserved under any global unitary transformation. These sets are convex and contain the maximally mixed state as well as states sufficiently close to it. Due to this invariant condition, AS states are completely determined from their spectrum.
@@
Our method provides novel simple analytical sufficient conditions for AS in arbitrary dimensions based on the knowledge of very few or even one eigenvalue of the spectrum. These analytical conditions correspond to convex simplexes that contain extremal points of the AS set, improving presently known conditions for AS. Moreover, we have demonstrated that AS conditions can be cast as convex functions of the form $f(\rho)\leq 0$. This fact allows us to unify all known AS conditions into a single and stronger one using convex programming. Such a technique can be generalized to any convex set to improve its characterization, which is normally a very hard task. Our method also yields new conditions on the problem of absolute separability in the multipartite scenario, as well as in the case of the symmetric subspace. Finally, we provide some new criteria on the twin problem of AP.
Let us conclude that the criteria presented in this work rely on the knowledge of the lowest eigenvalues of the state. In contrast, the majority of setups aimed at the evaluation of eigenvalues focus on the maximal one [48, 49]. However, our results motivate the study of the experimental setups that lower-bound the spectrum of a given state $\rho$, prompting new theoretical and experimental avenues on the problem of separability.
@@
Acknowledgements.– We thank A.Diebra and J.Romero-Pallejà for insightful comments and discussions with J. Ahiable, K.N.B. Teja, J.Tura and A. Winter.
@@
We acknowledge financial support from Spanish MICIN (projects: PID2022:141283NB-I00;139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda. We also acknowledge support from: Europea Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033), EU-QUANTera Projects: MAQS PCI2019-111828-2, DYNAMITE PCI2022-132919, ExTRaQT-PCI2022-132965. Support from Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024) funded by the European Union are acknowledged. This project has received funding from EU Horizon Europe Program ORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690; FET-OPEN OPTOlogic, Grant No 899794, NeQSTGrant-01080086. ICFO Internal “QuantumGaudi" project. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Derivation of a single linear inequality from the convex hull of two simplexes
In this appendix, we present the detailed procedure for deriving the inequalities described in Eqs. (6), (16), (18), (20), (22) and (24) which arise from the geometric structure of the two simplexes in the eigenvalue space, given by positivity of the inverse of the reduction map $\Lambda_{\alpha}^{-1}(\rho)\geq 0$ (3). For the low-dimensional specific cases, we also describe the procedure to check they become an AS condition, while giving numerical evidence for arbitrary dimensions that Eqs. 6 and 20 to also hold as sufficient conditions for AS, besides containing the whole CH of the two linearly independent simplexes (proven in Appendix B).
In the context of absolute separability (AS) of systems of total dimension $D$, sufficient conditions are inherently dependent on the spectrum $\bm{\lambda}$ of the quantum state $\rho$. Consequently, it is often more practical to work within the space of eigenvalues $\{\lambda_{i}\}$ rather than over the entire space of physical states.
From the positivity of the inverse of the reduction map (3), we derive $D$ inequalities for the $D$ eigenvalues of the state $\rho$ for each value of $\alpha$. Theorem 3 provides a simplified form of these inequalities, given a specific ordering of the eigenvalues. The $D$ conditions enclose a regular $(D-1)$-dimensional simplex, which center is in the maximally mixed state (MMS). The vertices of the simplex are given by all the different permutations of the vector $(1-(D-1)/(D+\alpha),1/(D+\alpha),\cdots,1/(D+\alpha))$.
The unit trace constraint allows us to eliminate a single variable, $\sum_{i}\lambda_{i}:=\tilde{\lambda}_{0}=1$. Next, the orthogonal complement of the vector $(1,\cdots,1)$ is used to complete a basis of the $D-1$ subspace describing the barycentric coordinates (see Figure 1 of the main text).
The convex hull (CH) formed by both simplexes represents a polytope whose vertices are the union of the vertex sets from each simplex. While the resulting polytope is not regular, it exhibits symmetry under eigenvalue permutations, reflecting the identical ordering possibilities of the eigenvalues. In computational geometry, the $H-$representation (half-space representation) of a polytope involves describing it as an intersection of half-spaces. These half-spaces can be represented by linear inequalities of the form $A\bm{\lambda}\leq\mathbf{b}$, being $A$ a matrix, but since the desired object is invariant under permutations, the matrix $A$ will be in this case a matrix of $D!$ identical rows (the possible $D!$ orderings of the eigenvalues [50]).
For low dimensional systems, it is possible to compute the set of linear inequalities analytically, by computing the hyperplanes given by each possible set of $D-1$ points, and considering the most external ones as the ones containing the vertices of the biggest dual polytope. Given the difficulty of extending the calculations for all $D$, a numerical approach has been used to compute the equations of the facets of the polyhedron. In specific, making use of the python library pypoman - Polyhedron manipulation in Python [51]. The direct output of the package is the matrix $A$ and the vector $\mathbf{b}$ for a given set of points (the vertex of both simplexes).
In $D-1$-dimensional space, these inequalities depend on the orientation of the basis orthogonal to the vector $(1,\cdots,1)$, not only on the dimension of the Hilbert space. To generalize these inequalities across different dimensional systems, we reexpress them in the full $D$-dimensional eigenvalue space. This allows for a more straightforward comparison between different quantum systems, and the resulting linear matrix inequalities take the form $A\bm{\lambda}\leq\mathbf{b}$, where $\mathbf{b}$ is a constant vector, and $A$ is a matrix of permuted vectors $\mathbf{c}=(p,\cdots,p,r,-q,\cdots,-q)$. Here, the multiplicities of $p$, $q$, and $r$ are functions of the system’s dimension, ensuring that the inequalities remain invariant under permutations. Being $m(p)$ the multiplicity of $p$ in the given vector, the size of the matrix is given by $D\times\binom{D}{\max{m(p),m(q)}}\leq D!$. Where $p,r,q,s$ can be expressed as a function of the total dimension of the system.
It is necessary to find the expression that corresponds to the ordered eigenvalues, i.e., the one that encloses the specific region in the polytope in non-decreasing order ($\mathbf{c}$). Reapplying the normalization condition $\sum_{i}\lambda_{i}=1$, one can eliminate the dependence on all the eigenvalues whose coefficient would be $p$ in $\mathbf{c}\cdot\bm{\lambda}^{T}$.
It is important to note that these expressions have not been derived in a general manner, but instead from a bottom-up approach, where we have computed the first low-dimensional cases and generalized the dependence of $p,q,r,s$ on the total dimension of the system, checking the validity for some high dimensional cases.
In general, it is possible to find in a case by case basis the $D-1$ points that generate the hyperplane, as well as to check that it encloses the bigger dual polytope. By considering the intersection of the external hyperplanes with the ordering, it is also possible to see that one single inequality is enough to establish the criteria. Nevertheless, for high dimensions, this procedure becomes intractable even though the python library computes the most external ones by construction.
An example of some gathered numerical evidence can be done by uniformly sampling random states. Then, the disciplined convex program based on Theorem 4:
$\begin{array}[]{crl}\min_{\{\varrho_{1},\varrho_{2}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\varrho_{1}+\varrho_{2}\\ &\lambda_{\min}(\varrho_{1})&\geq\frac{\mathrm{Tr}(\varrho_{1})}{D+\alpha_{+}}\\ &\lambda_{\max}(\varrho_{2})&\leq\frac{\mathrm{Tr}(\varrho_{2})}{D-|\alpha_{-}|}\\ \end{array},$ (25)
being $\alpha_{\pm}$ the extreme values of $\alpha$, identifies the same sets of states as those found through the linear inequalities derived from the eigenvalues.
Appendix B Proof that the convex hull satisfies Proposition 1
@@
After obtaining numerically the convex hull inequality, as described in Appendix A, we proceed to verify that the convex hull of two simplexes satisfies such inequality. To this end, we revisit key concepts of majorization, as outlined in [52], along with the references within the text.
@@
To begin with, if a vector $\mathbf{x}$ is majorized by $\mathbf{y}$, that is, $\mathbf{x}\prec\mathbf{y}$ then $\mathbf{x}$ is less spread than $\mathbf{y}$. In the context of the space of eigenvalues of quantum density matrices, we refer as $\mathbf{x}$ being more mixed than $\mathbf{y}$. Formally,
@@
From the previous lemma, one can immediately infer that any point in the space of eigenvalues of a density matrix inside the simplexes given by Eq. (3) is majorized by any of the extreme points of the simplex, which are all different permutations of a single vertex.
@@
Since tha majorization condition requiers the fulfillement of the normalization Eq. (26), the previous lemma implies that any point inside of the convex hull of the two simplexes is majorized by some points on the border of the convex hull (obtained as a linear combination of the vertex of the simplexes).
@@
We now verify that any vector contained in the convex hull of the two simplexes fulfills the proposed condition. First, given the eigenvalues of a density matrix in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,D-1}$ we define
$f(\bm{\lambda})=1-3\cdot\sum_{i=0}^{\lfloor\frac{D+1}{3}\rfloor-1}\lambda_{i}+\left(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor\right)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor},$ (28)
from Eq. (6) such that our condition is expressed as $f(\bm{\lambda})\leq 0$ (c.f. Section 3.3). Here, the argument of the function is any vector of eigenvalues, which is not necessarily sorted. This function is clearly symmetric under permutations, as it accounts for the ordering of each component of the vector independently of its position. Furthermore, it can be verified from Lemma 14 that for natural values of $D,N\in\mathbb{N}$, the function $f(\bm{\lambda})$ is Schur-convex. On the other hand, we recall that the vertices of the AS simplexes are given by all the permutations of the vectors $\bm{\lambda}_{\pm}=\left(\frac{1}{D+\alpha_{\pm}},\cdots,\frac{1}{D+\alpha_{\pm}},1-\frac{D-1}{D+\alpha_{\pm}}\right)$, with $\alpha_{+}=2,\alpha_{-}=-1$ in the space of eigenvalues. Due to the convexity of the AS set and Schur-convexity of the function, we check the inequality $f(\bm{\lambda})\leq 0$ on the extremals $\bm{\lambda}_{\pm}$. In this regard, we confirm the tight value $f(\bm{\lambda}_{\pm})=0$. Thus, it follows from Schur-convexity that all the vectors contained in the convex hull of the points generated by the permutations of $\bm{\lambda}_{\pm}$ fulfill the inequality. Finally, the procedure outlined here can be extended to Eqs. (16), (18), (20), (22) and (24) of the main text.

### v2

@@ 4.1 Symmetric absolute separability and absolute PPT
[…]
$\sigma^{T_{A}}=\bigoplus_{i\neq j}\begin{pmatrix}\frac{1}{3}&\alpha x_{i}x_{j}\\ \alpha x_{i}x_{j}&\frac{1}{3}\end{pmatrix}\bigoplus_{k}\underbrace{\begin{pmatrix}1+\alpha x_{k}^{2}&\frac{1}{3}&\cdots&\cdots&\cdots&\frac{1}{3}\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{6}&\cdots&\frac{1}{6}&\frac{1}{3}\\ \vdots&\frac{1}{6}&\frac{1}{3}&\cdots&\frac{1}{3}&\frac{1}{6}\\ \vdots&\vdots&&\ddots&&\vdots\\ \vdots&\frac{1}{6}&\frac{1}{3}&\cdots&\frac{1}{3}&\frac{1}{6}\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{6}&\cdots&\frac{1}{6}&\frac{1}{3}\end{pmatrix}}_{:=M_{k}}\bigoplus_{p\neq q\neq r}\begin{pmatrix}1/6&1/6\\ 1/6&1/6\end{pmatrix}$. (30)
Positivity of the first d(d1)/2d(d-1)/2 blocks implies $1/3-|\alpha|x_{d-1}x_{d-2}\geq 0$. From $x_{d-1}x_{d-2}\leq 1/2$, it yields the range $2/3\leq\alpha\leq 2/3$. On the other hand, for the next $d$ blocks, it is sufficient to check the positivity of $M_{d-1}$ with $x_{d-2}=0,x_{d-1}=1$. For $\alpha\geq 0$, $M_{d-1}\geq 0$. The minimal value of $\alpha$, $\alpha_{\min}<0$ such that $M_{d-1}\geq 0$ will decrease its rank. For this condition, we can construct a $d$-dimensional set of linearly independent vectors as the orthogonal complement of the matrix $M_{d-1}$ once the first row has been deleted. The first (d2)(d-2) ones are of the form (0,,0,1,0,,0,1,0,,0)(0,...,0,1,0,...,0,-1,0,...,0) where the $1$ and 1-1 are set in the positions $1+\kappa$ and $\kappa$ of the vector respectively, with $\kappa=1,2,..,d-2$. The ones restraining the value of alpha are of the form $(-d/2,1,...,1,0,...,0):=\mathbf{v}_{1}$ and $(-d/2,1,...,1,0,1,0,...,0):=\mathbf{v}_{2}$, both of them containing $d/2$ ones. It can be seen trivially that all the vectors of the first form are orthogonal to the first row $(1+\alpha,1/3,...,1/3):=\mathbf{w}$ of $M$. The equality $\mathbf{v}_{i}\cdot\mathbf{w}=0$ is fulfilled by $\alpha_{\min}=-(d+2)/(3d)$. Thus, for $\alpha\geq-(d+2)/(3d)$, the last block is PSD as well.
From the previous considerations, we conclude $\alpha_{-}=-(d+2)/(3d)$ and $\alpha_{+}=2/3$. Finally, Eq. (29) follows from Lemma 1 with the announced bounds on $\alpha$. ∎
Remarkably, we present analytical tight bounds of the parameter $\alpha_{\pm}$ for arbitrary dimensions $d$ and $N=2,3$. Contrarily to the bounds for $\alpha_{-}$ in Proposition 2 , as expected from the analytical derivations presented in Theorems 9 and 10, it is possible to find NPT states arbitrarily close to the given extreme values of $\alpha_{-}$. In general, these bounds are algebraic numbers defined by the roots of characteristic polynomials of high degree. For $N>3$, the structure of the matrices is more involved (see e.g. recent contribution [55]). The systematic extension of our results to arbitrary number of partitions $N$ warrants further investigation and it will be examined elsewhere.
@@ 5 Conclusions
In this work we have presented a new method to tackle the problem of absolute separability and absolute PPT based on inverting non-completely positive maps. The AS (AP) sets are defined as those formed by the states whose separability (positivity under partial transposition) is preserved under any global unitary transformation. These sets are convex and contain the maximally mixed state as well as states sufficiently close to it. Due to this invariant condition, AS and AP states are completely determined from their spectrum.
@@
Our method provides novel simple analytical sufficient conditions for AS in arbitrary dimensions based on the knowledge of very few or even one eigenvalue of the spectrum. These analytical conditions correspond to convex simplexes that contain extremal points of the AS set, improving presently known conditions for AS. Moreover, we have demonstrated that AS conditions can be cast as convex functions of the form $f(\rho)\leq 0$. This fact allows us to unify all known AS conditions into a single and stronger one using convex programming. Such a technique can be generalized to any convex set to improve its characterization, which is normally a very hard task. Our method also yields new conditions on the problem of absolute separability in the multipartite scenario, as well as in the case of the symmetric subspace. Finally, from the same methodology, we provide various new criteria on the twin problem of AP for multipartite symmetric systems.
While some protocols for teleportation [56] or error correction [57] aim at the evaluation of the maximal eigenvalue of the state, our criteria mostly rely on the knowledge of the lowest one, motivating the study of experimental setups that lower-bound the spectrum of a given state. Moreover, our results of absolute separability in the multipartite scenario highlights the present lack of results in characterizing the entanglement properties of highly mixed systems, prompting deeper studies on the topic.
Finally, it is worth noting that similar techniques as the one we report here can be used to study the invariance of other convex sets of states under global unitary transformations. For instance, our method can be applied to further analyze absolute non-violation of steering [58] or Bell-CHSH [59] inequalities. It can also be used to introduce and characterize other invariant convex sets, such as absolute Schmidt Number [60]. Lastly, the possibility of exploiting these absolute sets in a resource theory remains an open question.
@@
Acknowledgements. – We thank A. Diebra and J. Romero-Pallejà for insightful comments and acknowledge fruitful discussions with E. Serrrano-Ensátiga, J. Ahiable, K.N.B. Teja, J. Tura and A. Winter.
@@
J.A.-V. acknowledges financial support from Ministerio de Ciencia e Innovación of the Spanish Goverment FPU23/02761. We acknowledge financial support from Spanish MICIN (projects: PID2022: 141283NB100; 139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda. We also acknowledge support from: Europea Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033), EU-QUANTera Projects: MAQS PCI2019-111828-2, DYNAMITE PCI2022-132919, ExTRaQT-PCI2022-132965. Support from Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024) funded by the European Union are acknowledged. This project has received funding from EU Horizon Europe Program ORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690; FET-OPEN OPTOlogic, Grant No 899794, NeQSTGrant-01080086. ICFO Internal “QuantumGaudi" project. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. G. R.-M. acknowledges funding from the European Innovation Council accelerator grant COMFTQUA, no. 190183782.
Appendix A Numerical approach to Lemma 1
In this appendix, we present a numerical method to compute the linear inequalities defining the convex hull of a set of extremal points of the absolute separability (AS) set. This approach is particularly useful when the assumptions underlying Lemma 1 are not met, when the symmetry between the primal and dual simplices is broken, as it might be the case for other sets of extremal points [29, 50].
Working in the eigenvalue $\{\lambda_{i}\}$ space, normalized by the trace constraint $\sum_{i}\lambda_{i}=1$, allows the AS criteria to be expressed in terms of convex polytopes in (D1)(D-1)-dimensional space. A basis orthogonal to (1,,1)(1,...,1) provides a convenient coordinate system. The polytope can be represented in its $H$-representation as a set of inequalities $A\lambda\leq b$, where $A$ and $b$ are derived from the vertices of the convex hull. In cases with permutation invariance, $A$ consists of all permutations of a single inequality vector [61].
The analitical computation of $A,b$ is difficult to extend to high dimensions. Instead, numerical tools such as pypoman [62] can be used to compute $A$, $b$. Since in the normalized space $A,b$ might depend on the basis chosen, it is advisable to convert them into expressions of the whole space of eigenvalues. Since the representation might depend on the chosen basis, the inequalities are lifted back to the full $D$-dimensional eigenvalue space to preserve generality.
To identify a minimal set of necessary inequalities for a fixed eigenvalue ordering, we consider removing a candidate inequality $\mathbf{a}^{T}\bm{\lambda}\leq\beta$ and augmenting the remaining system with the normalization and the specific ordering ($\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,1,...,D-2}$) constraints. Denoting the new system as $A^{\prime}\bm{\lambda}\leq\mathbf{b}^{\prime}$, we solve the linear program
$\text{max}\{\mathbf{a}^{T}\bm{\lambda}|A^{\prime}\bm{\lambda}\leq\mathbf{b}^{\prime}\}=\gamma.$ (31)
If $\gamma\leq\beta$, the removed inequality is redundant. Repeating this process allows reduction to the minimal necessary set.
For the case of Theorem 3, the vertices $\mathbf{e}_{i}$ are permutations of $(1-(D-1)/(D+\alpha),1/(D+\alpha),...,1/(D+\alpha))$. The resulting inequalities in the full eigenvalue space take a form such that $\mathbf{b}$ is a constant vector, and $A$ is a matrix of permuted vectors $\mathbf{c}=(p,\cdots,p,r,-q,\cdots,-q)$. The coefficients $p,q,r$ and their multiplicities depend only on $D,\alpha_{\pm}$, ensuring permutation invariance. The number of inequalties scales as $D\times\binom{D}{\max{m(p),m(q)}}\leq D!$. Moreover, the reduction on the number of inequalities can be done until just one expression is needed a part from the ordering and the normalization expressions.
Appendix B Schur convexity
@@
In this section, we revisit key concepts of majorization, as outlined in [63], along with the references within the text.
@@
To begin with, if a vector $\mathbf{x}$ is majorized by $\mathbf{y}$, that is, $\mathbf{x}\prec\mathbf{y}$ then $\mathbf{x}$ is less spread than $\mathbf{y}$. In the context of the space of eigenvalues of quantum density matrices, we refer as $\mathbf{x}$ being more mixed than $\mathbf{y}$. Thus, majorization theory is closely related with absolute separability and PPT. Formally,
@@
From the previous lemma, one can immediately infer that any point in the space of eigenvalues of a density matrix inside the simplexes given by Eq. (3) is majorized by any of the vertices of the simplex, which are all different permutations of a single one. The same notion can be applied to any of the extreme points that might be derived for AS and AP for better characterization of both sets.
@@
Since the majorization condition requires the fulfillement of the normalization Eq. (32), the previous lemma implies that any point inside of the convex hull of the two simplexes is majorized by some points on the border of the convex hull (obtained as a linear combination of the vertex of the simplexes).
@@
We now verify that any vector contained in the convex hull of the two simplexes fulfills the proposed condition in Lemma 1. First, given the eigenvalues of a density matrix in non-decreasing order $\{\lambda_{k}\leq\lambda_{k+1}\}_{k=0,\cdots,D-2}$ we define
$f(\bm{\lambda})=1-\left(1-\frac{\alpha_{+}}{\alpha_{-}}\right)\cdot\sum_{i=0}^{c-1}\lambda_{i}-\left[D+\alpha_{+}+c\cdot\left(\frac{\alpha_{+}}{\alpha_{-}}-1\right)\right]\cdot\lambda_{c},$ (34)
from Eq.(8) such that our condition is expressed as $f(\bm{\lambda})\leq 0$ (cf. Section3.3). Here, the argument of the function is any vector of eigenvalues, which is not necessarily sorted. This function is clearly symmetric under permutations, as it accounts for the ordering of each component of the vector independently of its position. Furthermore, it can be verified from Lemma 5 that for natural values of $D\in\mathbb{N}$, $\alpha_{+}>0$ and $\alpha_{-}<0$; the function $f(\bm{\lambda})$ is Schur-convex. On the other hand, we recall that the vertices of the AS simplexes are given by all the permutations of the vectors $\bm{\lambda}_{\pm}=\left(\frac{1}{D+\alpha_{\pm}},\cdots,\frac{1}{D+\alpha_{\pm}},1-\frac{D-1}{D+\alpha_{\pm}}\right)$, in the space of eigenvalues. Due to the convexity of the AS set and Schur-convexity of the function, we check the inequality $f(\bm{\lambda})\leq 0$ on the extremals $\bm{\lambda}_{\pm}$. In this regard, we confirm the tight value $f(\bm{\lambda}_{\pm})=0$. Thus, it follows from Schur-convexity that all the vectors contained in the convex hull of the points generated by the permutations of $\bm{\lambda}_{\pm}$ fulfill the inequality. Finally, the procedure outlined here can be extended to Eqs. (9), (19), (21), (23), (26) and (29) of the main text.